OpenStudy (anonymous):

if (x-3) is a factor of x^2-px-x, find the value of p

OpenStudy (loser66):

take long division or synthetic to get value of p, I got p = 2. How about you??

OpenStudy (anonymous):

same here. but my book says the ans is 1. I guess its wrong.

OpenStudy (loser66):

ok, let check: if p =1 , then the expression is \(x^2 -x -x = x^2-2x\) factor x out, we have x(x-2) , so that x -3 can't be its factor --> p=1 is not the right answer. Now, our result is p =2 then, the expression is \(x^2 -2x -x = x^2 -3x\) factor x out, we get x(x-3) . It shows x-3 is one of the factor of the expression \( x^2 -px -x\) when p =2. Proof done.